# A multi-stage Bayesian approach to fit spatial point process models

### v1

@@ Abstract
Spatial point process (SPP) models are commonly used to analyze point pattern data, including presence-only data in ecology. Current methods for fitting these models are computationally expensive because they require numerical quadrature and algorithm supervision (i.e., tuning) in the Bayesian setting. We propose a flexible and efficient multi-stage recursive Bayesian approach to fitting SPP models that leverages parallel computing resources to estimate point process model coefficients and derived quantities. We show how this method can be extended to study designs with compact observation windows and allows for posterior prediction of total abundance and points in unobserved areas, which can be used for downstream analyses. We demonstrate this approach using a simulation study and analyze data from aerial imagery surveys to improve our understanding of spatially explicit abundance of harbor seals (Phoca vitulina) in Johns Hopkins Inlet, a protected tidewater glacial fjord in Glacier Bay National Park, Alaska.
@@ 1 Introduction
Spatial point process (SPP) models are a class of generative stochastic models that give rise to a random number of irregularly spaced events in a spatial domain. They have been used to analyze point patterns in various fields, such as ecology (e.g., Haase, 1995; Law, 2009; Warton and Shepherd, 2010; Renner et al., 2015), seismology (e.g., Ogata, 1998), and social sciences (e.g., Mohler et al., 2011). One common approach to SPP modeling is to use spatially referenced covariates to fit an inhomogeneous Poisson process (IPP), a point process model in which point intensity varies across space (Gelfand et al., 2010). This approach is especially attractive in applications where point intensity is believed to depend on spatially varying environmental factors.
@@
The stochasticity in an SPP arises from two sources: the random number of point events and their locations, which we refer to as $n$ and $\{\mathbf{s}_{i}\}_{i=1}^{n}$, respectively. A standard IPP likelihood models both data types jointly, allowing spatial heterogeneity and total abundance (i.e., the total number of points in the study domain) to be modeled in a unified framework. This feature is especially useful when the data are partially observed (i.e., total abundance is unknown and may exceed $n$) and inference on total abundance is desired. We extend the IPP likelihood to data observed in compact observation windows, a special case of partially observed data defined by Baddeley et al. (2015). This may arise naturally from certain study designs, such as aerial imagery surveys.
@@
IPP likelihoods are often intensive to compute because they involve an integral over the spatial domain, which is analytically intractable and must be approximated using methods such as numerical quadrature. This is especially inconvenient for Monte Carlo Markov Chain (MCMC; Gelfand and Smith, 1990) methods because the integral is necessarily approximated for each iteration. Furthermore, conventional MCMC algorithms for fitting IPPs often require algorithmic tuning due to a lack of conjugate priors.
@@
One alternative to fitting a Bayesian model using conventional MCMC is to use recursive Bayes approaches, also referred to as Bayesian filtering or sequential inference, to fit the model in stages (Särkkä, 2013). Although these approaches are widely applicable, recursive Bayes approaches for fitting various ecological models have been developed in recent years due to their natural application to hierarchical models and computational efficiency (e.g., Hooten et al., 2023, 2024; Johnson et al., 2022; Leach et al., 2022; McCaslin et al., 2021). These methods often implement a version of the prior-proposal recursive Bayes (PPRB) method described by Hooten et al. (2021). In a two-stage implementation of PPRB, the data are partitioned into two subsets, $\mathbf{y}:=(\mathbf{y}_{1},\mathbf{y}_{2})^{\prime}$. We let $\boldsymbol{\theta}$ represent a vector of model parameters and $\boldsymbol{\theta}]$ represent its prior distribution. We use the bracket notation $\cdot$]’ to represent probability distributions hereafter (Gelfand and Smith, 1990). In the first stage, we use $\mathbf{y}_{1}$ to find a transient posterior distribution $\boldsymbol{\theta}|\mathbf{y}_{1}]\propto[\mathbf{y}_{1}|\boldsymbol{\theta}][\boldsymbol{\theta}]$. Next, in the second stage, we use the transient posterior distribution as a prior and proposal distribution in an MCMC algorithm to obtain a sample from the full posterior distribution
@@
Conventionally, the partitions comprise of data of the same type, but alternative perspectives on partitioning can be helpful in some cases. For instance, PPRB approaches for fitting capture-recapture models partition the data such that the two partitions comprise of the observed binary capture histories and the number of observed individuals, respectively (Hooten et al., 2023, 2024). Similarly, we show that it is natural to partition spatial point pattern data such that the first partition comprises the observed locations $\{\mathbf{s}_{i}\}_{i=1}^{n}$ and the second partition comprises the number of observed point events $n$. We can then fit the data model conditioned on $n$ in the first stage and assimilate $n$ in the second stage. This partitioning scheme allows us to take advantage of well-known strategies that use the conditional data model $\{\mathbf{s}_{i}\}|\boldsymbol{\theta},n]$, such as the logistic and Poisson regression approximations (Aarts et al., 2012; Baddeley et al., 2010; Warton and Shepherd, 2010). We propose a multi-stage method for fitting the IPP using PPRB, along with a variety of strategies for first-stage sampling from the transient posterior distribution, resulting in an efficient and flexible approach to fitting a Bayesian SPP model.
In section 2, we provide an overview for fitting an IPP with spatially explicit covariates. Next, in section 3, we construct a Bayesian IPP model, propose a PPRB procedure for fitting the model, and compare various first-stage sampling strategies. We then demonstrate the multi-stage approach on simulated data in a compact window setting in section 4. Finally, in section 5, we apply our approach to data from aerial imagery surveys to learn spatially explicit abundance of harbor seals in Johns Hopkins Inlet, a protected glacial fjord in Glacier Bay National Park, Alaska.
@@ 2 Inhomogeneous Poisson Process
We consider a space of interest, $\mathcal{S}\subset\mathbb{R}^{2}$, and a grid of disjoint subspaces $\{A_{l}\}_{l=1}^{L}$ such that $\mathcal{S}=\cup_{l}A_{l}$. We let $\mathbf{s}_{i}\in\mathcal{S}$ be a two-dimensional vector denoting event location for $i=1,\dots,n$ observed events. We also denote $N(\cdot)$ as a random counting measure and $N$ as the total number of events in $\mathcal{S}$ (i.e., $N=N(\mathcal{S})).$
@@
An IPP can also be thought of as the continuous limit of an independent Poisson count model for an increasingly fine grid of subspaces (i.e., as $L$ approaches $\infty$) over the spatial domain (Fithian and Hastie, 2013). Our method focuses on modeling the continuous intensity function $\lambda$ parametrically using a computationally efficient and flexible Bayesian approach. Additional details and theoretical properties on the IPP are available in Gelfand et al. (2010).
@@ 2.1 Complete Likelihood
Assuming we have $p$ covariates associated with each point in $\mathcal{S}$, we can model the IPP intensity as a function of the spatially referenced covariates $\lambda(\mathbf{s})=f(\beta_{0}+\mathbf{x}^{\prime}(\mathbf{s})\boldsymbol{\beta}\bigr{)}$, where $\beta_{0}$ represents the intercept, $\mathbf{x}$ and $\boldsymbol{\beta}$ denote $p\times 1$ vectors of covariates and their corresponding coefficients, respectively, and the function $f$ represents a positive strictly increasing differentiable function. We assume the convention $f(x)=\exp(x)$. If desired, this assumption may be relaxed by using a different link function and/or transforming the covariates using a basis expansion.
The IPP likelihood is then as follows
@@
for $i=1,\dots,n$. We refer to this IPP likelihood that jointly models $\{\mathbf{s}_{i}\}_{i=1}^{n}$ and $n$ as the complete likelihood hereafter.
In practice, the integral $\Lambda(\mathcal{S})$ is not analytically tractable and must be approximated. Using covariate information for grid centers $\mathbf{u}_{l}$ for $l=1,\dots,L$, we can approximate $\Lambda(\mathcal{S})$ with numerical quadrature and use $\{\mathbf{u}_{l}\}_{l=1}^{L}$ as quadrature points, but this can be computationally intensive, especially for large $L$.
@@ 2.2 Conditional Likelihood
for $i=1,\dots,n$, where $\lambda(\mathbf{s})=\exp(\mathbf{x}^{\prime}(\mathbf{s})\boldsymbol{\beta}\bigr{)}$. We refer to this as the conditional likelihood hereafter. It is important to note that unlike in the complete likelihood, $\beta_{0}$ is non-identifiable because it cancels in the numerator and denominator in (2).
The conditional likelihood, sometimes referred to as the conditional IPP, is used in the first stage of our multi-stage algorithm. Additionally, it is used in applications where the number of observed points $n$ is treated as known and fixed a priori. For instance, in telemetric surveys for studying animal movement (Hooten et al., 2017), it is common to fit an SPP model to the observed spatially explicit points for a predetermined number of animal relocations (e.g., Johnson et al., 2013). The conditional likelihood suffers from similar computational challenges in approximating $\Lambda(\mathcal{S})$; however, useful likelihood approximations have been developed to circumvent this issue.
@@ 2.2.1 Logistic Regression Approximation
One widely adopted approximation for the conditional likelihood in (2) is the logistic regression approximation, which involves augmenting $m$ background points to the data to represent locations where point events were absent (Baddeley et al., 2010; Warton and Shepherd, 2010; Fithian and Hastie, 2013). Warton and Shepherd (2010) showed that if the number of observed points $n$ is fixed and the number of background points $m$ grows infinitely large, then the logistic regression coefficients (intercept excluded) converge to the maximum likelihood estimates of the IPP coefficients in the conditional likelihood (2).
To implement this approach, we use the Berman-Turner device (Berman and Turner, 1992), where we define an indicator variable, $y(\mathbf{s})=\mathds{1}(\mathbf{s}\in\{\mathbf{s}_{i}\}_{i=1}^{n})$, which indicates whether a location $\mathbf{s}$ is within our set of observed points. By definition, we have $y(\mathbf{s}_{i})=1$ for $i=1,\dots,n$. To obtain points where $y(\mathbf{s})=0$, we generate a uniform random sample of size $m$ within $\mathcal{S}$. For guidance on choosing the number of background points $m$ in practice, see Northrup et al. (2013).
For $m$ background points, we order $y(\mathbf{s}_{i})$ such that $y(\mathbf{s}_{i})=1$ for $i=1,\dots,n$ and $y(\mathbf{s}_{i})=0$ for $i=n+1,\dots,\tilde{n}$ where $\tilde{n}=n+m$. We then fit $y(\mathbf{s}_{i})\sim\text{Bern}(p_{i})$ where $\text{logit}(p_{i})=\beta_{0}+\mathbf{x}^{\prime}(\mathbf{s}_{i})\boldsymbol{\beta}.$ Although the logistic regression intercept is not relevant for the IPP model, note that its inclusion is necessary to adjust for the background sample size (Fieberg et al., 2021).
@@ 2.3 Compact Observation Window Extension
for $i=1,\dots,n$ and $j=1,\dots,J$ where $n=\sum_{j}n_{j}$. We refer to (3) as the windowed complete likelihood hereafter. The windowed conditional likelihood can be found by conditioning on $\{n_{j}\}$.
@@ 3 Bayesian Model
Using the complete likelihood in (1), we can estimate and quantify uncertainty for the parameters of interest, $\beta_{0}$ and $\boldsymbol{\beta}$, using a Bayesian model. The full posterior distribution is as follows:
@@
Recall that, for the complete likelihood in (1), $\Lambda(\mathcal{S})$ is analytically intractable. Using numerical quadrature to approximate this integral often involves computing a large sum that scales with the number of quadrature points. Furthermore, the integral must be computed at each MCMC iteration.
Secondly, we are unable to obtain closed form representations for the full-conditional distributions $\beta_{0}|\cdot]$ and $\boldsymbol{\beta}|\cdot]$. If we use a normal random walk proposal for $\beta_{0}$ and $\boldsymbol{\beta}$, we must perform tuning, which incurs additional computational costs.
Instead of evaluating the full posterior distribution in (4) using a conventional single-stage MCMC algorithm, we propose a multi-stage MCMC algorithm for fitting the complete likelihood that does not require parameter tuning and allows for numerical integration to be computed in parallel.

### v2

@@ Abstract
Spatial point process (SPP) models are commonly used to analyze point pattern data in many fields, including presence-only data in ecology. Existing exact Bayesian methods for fitting these models are computationally expensive because they require approximating an intractable integral each time parameters are updated and often involve algorithm supervision (i.e., tuning in the Bayesian setting). We propose a flexible, efficient, and exact multi-stage recursive Bayesian approach to fitting SPP models that leverages parallel computing resources to obtain realizations from the joint posterior, which can then be used to obtain inference on derived quantities. We outline potential extensions, including a framework for analyzing study designs with compact observation windows and a neural network basis expansion for increased model flexibility. We demonstrate this approach and its extensions using a simulation study and analyze data from aerial imagery surveys to improve our understanding of spatially explicit abundance of harbor seal (Phoca vitulina) pups in Johns Hopkins Inlet, a protected tidewater glacial fjord in Glacier Bay National Park, Alaska.
@@ 1 Introduction
Spatial point process (SPP) models are a class of generative stochastic models that give rise to a random number of irregularly spaced events in a spatial domain. They have been used to analyze point patterns in various fields, such as ecology (e.g., Haase, 1995; Law, 2009; Warton and Shepherd, 2010; Renner et al., 2015), seismology (e.g., Ogata, 1998), and social sciences (e.g., Mohler et al., 2011). Many models have been developed for SPP data including, but not limited to, the homogeneous and inhomogeneous Poisson point process, the log-Gaussian Cox process (LGCP; Møller et al., 1998), and the Hawkes process (Hawkes, 1971). This article focuses on a computational method for fitting the inhomogeneous Poisson Process (IPP), a point process model in which point intensity varies across space, and extensions that allow for increased model flexibility.
@@
The stochasticity in SPP data arises from two sources: the random number of point events and their locations, which we refer to as $n$ and $\{\mathbf{s}_{i}\}_{i=1}^{n}$, respectively. Although it is common in the species distribution modeling (SDM) literature to fit an IPP model using a likelihood conditioned on $n$ (e.g., Johnson et al., 2013), we focus on a complete IPP likelihood that models both data types jointly, allowing spatial heterogeneity and total abundance (i.e., the total number of points in the study domain) to be modeled in a unified framework. This feature is especially useful when the data are partially observed (i.e., total abundance is unknown and may exceed $n$) and inference on total abundance and/or prediction in unobserved areas is desired. We extend the complete IPP likelihood to data observed in compact observation windows, a special case of partially observed data defined by Baddeley et al. (2015) that may arise naturally from certain study designs, such as aerial imagery surveys.
@@
IPP likelihoods are often intensive to compute because they involve an integral over the spatial domain, which is analytically intractable and must be approximated using methods such as numerical quadrature. This is especially inefficient when using Monte Carlo Markov Chain (MCMC; Gelfand and Smith, 1990) methods because the integral is necessarily approximated each time the parameters are updated. Furthermore, conventional MCMC algorithms for fitting IPP models often require algorithmic tuning due to a lack of conjugate prior distributions. Various fast approximate Bayesian methods for fitting SPP models, such as the Integrated Nested Laplace Approximation (INLA; Rue and Martino, 2009) approach to fitting LGCP models (Illian et al., 2012), circumvent this issue by approximating marginal posterior distributions. However, obtaining exact Bayesian inference for the joint posterior distribution remains computationally challenging. Obtaining realizations from the joint posterior distribution, as opposed to approximating marginal distributions, may be beneficial for estimating derived quantities of interest, such as total abundance.
@@
One alternative to fitting an exact Bayesian model using conventional MCMC is to use recursive Bayes approaches, also referred to as Bayesian filtering or sequential inference, to fit the model in stages (Särkkä2013). Although these approaches are widely applicable, recursive Bayes approaches for fitting various ecological models have been developed in recent years due to their natural application to hierarchical models and computational efficiency (e.g., Hooten et al., 2023, 2024; Johnson et al., 2022; Leach et al., 2022; McCaslin et al., 2021). These methods often implement a version of the prior-proposal recursive Bayes (PPRB) method described by Hooten et al. (2021). In a two-stage implementation of PPRB, the data are partitioned into two subsets, $\mathbf{y}:=(\mathbf{y}_{1},\mathbf{y}_{2})^{\prime}$. We let $\boldsymbol{\theta}$ represent a vector of model parameters and $\boldsymbol{\theta}]$ represent its prior distribution. We use the bracket notation $\cdot$]’ to represent probability distributions hereafter (Gelfand and Smith, 1990). In the first stage, we use $\mathbf{y}_{1}$ to find a transient posterior distribution $\boldsymbol{\theta}|\mathbf{y}_{1}]\propto[\mathbf{y}_{1}|\boldsymbol{\theta}][\boldsymbol{\theta}]$. Next, in the second stage, we use the transient posterior distribution as a prior and proposal distribution in an MCMC algorithm to obtain a sample from the full posterior distribution
@@
Conventionally, the partitions in PPRB comprise data of the same type, but alternative perspectives on partitioning can be helpful in some cases. For instance, PPRB approaches for fitting capture-recapture models partition the data such that the first partition comprises the observed binary capture histories and the second partition comprises the number of observed individuals (Hooten et al., 2023, 2024). Similarly, we show that it is natural to partition spatial point pattern data such that the first partition comprises the observed locations $\{\mathbf{s}_{i}\}_{i=1}^{n}$ and the second partition comprises the number of observed point events $n$. We can then fit the data model conditioned on $n$ in the first stage and assimilate $n$ in the second stage. This partitioning scheme allows us to take advantage of well-known strategies that use the conditional data model $\{\mathbf{s}_{i}\}|\boldsymbol{\theta},n]$, such as the logistic and Poisson regression approximations (Aarts et al., 2012; Baddeley et al., 2010; Warton and Shepherd, 2010). We propose a multi-stage MCMC method for fitting the IPP model using PPRB, along with a variety of strategies for first-stage sampling from the transient posterior distribution, resulting in an efficient and flexible approach to fitting an exact Bayesian SPP model.
In section 2, we provide an overview for fitting an IPP model and potential extensions for additional model flexibility. Next, in section 3, we construct a Bayesian IPP model, propose a recursive Bayes procedure for fitting the model, and compare various first-stage sampling strategies. We then demonstrate the multi-stage approach on simulated data in a compact window setting in section 4. Finally, in section 5, we apply our approach to data from aerial imagery surveys to learn spatially explicit abundance of harbor seals in Johns Hopkins Inlet, a protected glacial fjord in Glacier Bay National Park, Alaska, and describe strategies to account for additional spatial heterogeneity in the model.
@@ 2 Inhomogeneous Poisson Process
We consider a space of interest, $\mathcal{S}\subset\mathbb{R}^{2}$, and a set of disjoint subspaces $\{A_{l}\}_{l=1}^{L}$ such that $\mathcal{S}=\cup_{l}A_{l}$. We let $\mathbf{s}_{i}\in\mathcal{S}$ be a two-dimensional vector denoting event location for $i=1,\dots,n$ observed events. We also denote $N(\cdot)$ as a random counting measure and $N$ as the total number of events in $\mathcal{S}$ (i.e., $N=N(\mathcal{S})).$
@@
An IPP can also be interpreted as the continuous limit of an independent Poisson count model for an increasingly fine grid of subspaces (i.e., as $L$ approaches $\infty$) over the spatial domain (Fithian and Hastie, 2013). Additional details and theoretical properties on the IPP are available in Illian et al. (2008) and Gelfand et al. (2010).
For $p$ covariates associated with each point in $\mathcal{S}$, we can model the IPP intensity $\lambda$ as a function of the spatially referenced covariates $\lambda(\mathbf{s})=f(\beta_{0}+\mathbf{x}^{\prime}(\mathbf{s})\boldsymbol{\beta}\bigr)$, where $\beta_{0}$ represents the intercept, $\mathbf{x}$ and $\boldsymbol{\beta}$ denote $p\times 1$ vectors of covariates and their corresponding coefficients, respectively, and the function $f$ represents a positive strictly increasing differentiable function. We assume the convention $f(x)=\exp(x)$. If desired, this assumption may be relaxed by using a different link function and/or transforming the covariates using a basis expansion. We discuss the latter in section 5.
@@ 2.1 Complete Likelihood
The IPP likelihood that jointly models the number of observed points and their locations is as follows
@@
for $i=1,\dots,n$. We refer to this IPP likelihood in (1) as the complete likelihood hereafter.
In practice, the integral $\Lambda(\mathcal{S})$ in (1) is not analytically tractable and must be approximated. Using covariate information for grid centers $\mathbf{u}_{l}$ for $l=1,\dots,L$, we can approximate $\Lambda(\mathcal{S})$ with numerical quadrature and use $\{\mathbf{u}_{l}\}_{l=1}^{L}$ as quadrature points, but this can be computationally intensive, especially for large $L$.
@@ 2.2 Conditional Likelihood
for $i=1,\dots,n$. We refer to this as the conditional likelihood hereafter. It is important to note that unlike in the complete likelihood, $\beta_{0}$ is non-identifiable because it cancels in the numerator and denominator in (2); however, inference on $\beta_{0}$ is necessary to estimate total abundance for partially observed domains. We let $\boldsymbol{\beta}$ denote the vector of coefficient values excluding the intercept hereafter for clarity.
The conditional likelihood (2) is commonly used in SDM literature and is appropriate for applications where the number of observed points $n$ is treated as known and fixed a priori. For instance, in telemetric surveys for studying animal movement (Hooten et al., 2017), it is common to fit an SPP model to the observed spatially explicit points for a predetermined number of animal relocations (e.g., Johnson et al., 2013). The conditional likelihood suffers from similar computational challenges in approximating $\Lambda(\mathcal{S})$; however, useful likelihood approximations have been developed to circumvent this issue.
@@ 2.2.1 Logistic Regression Approximation
One widely adopted approximation for the conditional likelihood (2) in both point process and SDM literature is the logistic regression approximation, which involves augmenting $m$ background points, also referred to as pseudo-absences, to the data to represent locations where point events were absent (Baddeley et al., 2010; Warton and Shepherd, 2010; Fithian and Hastie, 2013). Warton and Shepherd (2010) showed that if the number of observed points $n$ is fixed and the number of background points $m$ grows infinitely large, then the logistic regression coefficients (intercept excluded) converge to the maximum likelihood estimates of the IPP coefficients in the conditional likelihood (2).
To implement this approach, we use the Berman-Turner device (Berman and Turner, 1992), where we define an indicator variable, $y(\mathbf{s})=\mathds{1}(\mathbf{s}\in\{\mathbf{s}_{i}\}_{i=1}^{n})$, which indicates whether a location $\mathbf{s}$ is within our set of observed points. By definition, we have $y(\mathbf{s}_{i})=1$ for $i=1,\dots,n$. To obtain points where $y(\mathbf{s})=0$, we generate a uniform random sample of size $m$ within $\mathcal{S}$ (see Northrup et al. (2013) for guidance on choosing the number of background points $m$ in practice). For $m$ background points, we order $y(\mathbf{s}_{i})$ such that $y(\mathbf{s}_{i})=1$ for $i=1,\dots,n$ and $y(\mathbf{s}_{i})=0$ for $i=n+1,\dots,\tilde{n}$ where $\tilde{n}=n+m$. We then fit $y(\mathbf{s}_{i})\sim\text{Bern}(p_{i})$ where $\text{logit}(p_{i})=\beta_{0}+\mathbf{x}^{\prime}(\mathbf{s}_{i})\boldsymbol{\beta}.$ Although the logistic regression intercept is not relevant for the IPP model, note that its inclusion is necessary to adjust for the background sample size (Fieberg et al., 2021).
@@ 2.3 Compact Observation Window Extension
for $i=1,\dots,n$ and $j=1,\dots,J$ where $n=\sum_{j}n_{j}$. We refer to (3) as the windowed complete likelihood hereafter. Similar to the conditional likelihood (2), the windowed conditional likelihood can be derived by conditioning on $\{n_{j}\}.$
@@ 3 Bayesian Model
Using the complete likelihood (1), we can estimate and quantify uncertainty for the parameters of interest, $\beta_{0}$ and $\boldsymbol{\beta}$, using a Bayesian model. The full posterior distribution associated with our model is as follows:
@@
Recall that, for the complete likelihood (1), $\Lambda(\mathcal{S})$ is analytically intractable. Using numerical quadrature to approximate this integral often involves computing a large sum that scales with the number of quadrature points. Furthermore, because $\Lambda(\mathcal{S})$ depends on $\beta_{0}$ and $\boldsymbol{\beta}$, the integral must be approximated each time the parameters are updated.
Secondly, we are unable to obtain closed form representations for the full-conditional distributions for $\beta_{0}$ and $\boldsymbol{\beta}$. If we use a normal random walk proposal for $\beta_{0}$ and $\boldsymbol{\beta}$, we must perform tuning, which incurs additional computational costs. Unsupervised MCMC algorithms, such as Hamiltonian Monte Carlo (HMC; Neal, 2011) with the No-U-Turn Sampler (NUTS; Hoffman and Gelman, 2014), may also be implemented, but require overhead calculations.
Instead of evaluating the full posterior distribution (4) using a conventional single-stage MCMC algorithm, we propose a multi-stage MCMC algorithm for fitting the model using the complete likelihood that does not require parameter tuning and allows for the $\Lambda(\mathcal{S})$ approximation to be computed in parallel.